A particle is moving in a circle of radius 4 cm with constant speed of 1cm/s find average speed, average velocity and average acceleration in a time interval from t=0 to t=T/4 where T is the time period of the particle .Give only their magnitudes.

Grade:11

2 Answers

Lavina singh
6 years ago
avg speed =tot distance/tot time
tot distance=2πr/4=2π cm
tot time=T/4 s
avg speed=2π/T/4=8π/T cm/s
avg velocity= tot displacement/tot time
tot displ=2r=42 cm
tot time=T/4 s
avg vel=4√2/T /4=162/Tcm/s
the avg acc is zero(in magnitude) since speed is constant.
Ankitnavyadav
4 years ago
A body is moving with constant speed 10m/sec along a circle of radius 14m.Find the average velocity of the body from t=0 to t=4.4 sec